# Exploring the Nostr Ecosystem: A Study of Decentralization and Resilience

### v1

@@ 5. Post Availability & Replication
In §4.3, we discovered that certain relays (including the larger ones hosting over 1 million posts) experience periods of downtime. However, one of the key design innovations of Nostr is that the posts can be replicated across multiple relays, so that the availability of posts will not be significantly impacted by the downtime of a few relays. Indeed, in §4.1 and §4.2, we found that many posts are replicated across multiple relays, ASes, and countries. Therefore, this section delves deeper into the matter. Based on our real-world operational dataset, we aim to evaluate the effectiveness of this replication design innovation in enhancing posts availability and the overhead associated.
@@ 5.1. Number of Post Replications
We first assess the replication of posts, at a per-post level. Figure 6 presents the CDF of the number of relays and ASes in which a post is replicated. We see remarkable levels of replication: the majority of posts are replicated across multiple relays and ASes, with the average being 34.6 and 12.4, respectively. While the high volume of post replications contributes to increased post availability, it also lead to an excess of redundancies. This could incur unnecessary costs, such as those associated with relay storage space. We delve deeper into this aspect in the subsequent paragraphs.
5.2. Post Availability
@@
To evaluate the availability of posts, we simulate relay failures, using our dataset. We sequentially remove the most pivotal relays and ASes, and then measure the number of posts that remain available. Note, we define a post as available if it is replicated on at least one relay. We employ two ways of ranking the most pivotal relays and ASes to remove: measured by the largest number of (i) posts and (ii) users. We perform an additional ranking for ASes based on the number of relays.
Figure 7a illustrates the percentage of available posts subsequent to the removal of the top X relays. In stark contrast to the Fediverse (Raman et al., 2019), Nostr’s availability remains impressively robust, hovering around 90% even after the top 50 relays, ranked by the number of users, are removed. However, when removing the top relays based on the number of posts, the availability remains above 90% only until the top 30th relay is removed, after which it experiences a sharp decline, culminating at 71% at the top 50th. That said, we argue that the availability is still satisfactory.
Figure 7b displays the proportion of accessible posts following the elimination of the top X ASes. Despite the failure of entire ASes, over 80% of posts remaining available even after the removal of the top 10 ASes. To provide a comparison, the post availability plunges to less than 10% after the removal of the top 10 ASes hosting the instances (Raman et al., 2019). This is largely driven by the more uniform distribution of relays across distinct ASes, making them more resilient to individual AS failures. These findings suggest a high post availability within the Nostr ecosystem, a result largely attributable to the extensive replication of posts.
5.3. Post Redundancy
Despite the impressive post availability attained by Nostr, we suspect that this comes at the cost of excessive overhead. This involves an “over-replication” of some posts across numerous (unnecessary) relays. For example, our data shows that, 50% of posts are replicated across over 14 relays, and 25% are replicated across more than 49 relays. Arguably, this level of redundancy is unnecessary, except in the face of catastrophic failures. Consequently, we next measure the possible optimization space of post replications using our dataset.
The results are depicted in Figure 7c. By restricting the maximum replication to 20, we see that it is possible to eliminate 380 million post replications. Further reducing this limit to 10 leads to the removal of an additional 100 million post replications, totaling 480 million. This substantial reduction suggests that, if each user only posts to a maximum of 10-20 relays, over 400 million client-relay communications for posting could be saved, along with the storage space required for the over 400 million replications.
@@ 5.4. Duplicated Retrieval From Relays
The prior subsection has highlighted that Nostr’s approach to post replication substantially improves availability, but at the cost of storage overhead. This approach also creates another unexpected overhead: by connecting to multiple relays, consumers will redundantly retrieve the same post from several locations, resulting in unnecessary bandwidth usage. This occurs because consumers maintain connection to multiple relays, and retrieve any post of their following users from all locations. For example, if user $A$ follows user $B$, and both are using relays $R_{1}$, $R_{2}$, and $R_{3}$, then $B$ will send their post to all three relays. Consequently, $A$ will receive $B$’s identical post from $R_{1}$, $R_{2}$, and $R_{3}$ respectively. However, the ideal scenario would be for $A$ to retrieve the post from just one of the three relays. To date, Nostr has no solution to elegantly deal with this duplicated retrieval. We further explore this issue in this subsection.
5.4.1. Estimating traffic volume
@@
We begin by estimating the amount of wasted traffic caused by the duplicated retrieval of posts. Here, we assume that a user will retrieve all the posts from the users they are following via all the relays they are using, as defined by the follower relationships. First, we construct a directed graph $G$ of all users to represent following relationship. In this graph, a vertex $V$ signifies a user, and a directed edge $(U,V)$ indicates that user $U$ follows user $V$.
@@
Retrieval Count of a Post over an Edge.  For a post, $P$, that is shared over a following edge $(U,V)$, we calculate the number of times it is retrieved across relays. We refer to this as the retrieval count. First, we get $S_{P}^{R}$, which represents the set of relays where the post $P$ is duplicated. Second, we get $S_{U}^{R}$, which represents the set of relays utilized by user $U$. Last, we compute $|S_{P}^{R}\cap S_{U}^{R}|$, which corresponds to the retrieval count of $P$.
@@
Traffic over an Edge.  For a following edge $(U,V)$, we then calculate the volume of traffic over it. First, we get $S_{V}^{P}$, which represents all posts published by user $V$. Note, these are equal to all posts transmitted over edge $(U,V)$. Second, for every post in $S_{V}^{P}$, we calculate its retrieval count. Subsequently we then calculate the retrieval size, by multiplying the retrieval count by the size of the post. Last, we sum the retrieval counts and retrieval traffic for all posts within $S_{V}^{P}$ to obtain the total retrieval count and retrieval size over the edge $(U,V)$.
@@
Traffic over the Following Graph.  We also calculate the volume of traffic over the entire follower graph $G$. For every edge in $G$, we compute the traffic over the edge, and sum up the results. The pseudo code of the above process is listed in Algorithm 1.
5.4.2. Results
We use the above method to estimate the quantified amount of traffic. In total, the follower graph $G$ contains 3.5M follower relationships between users. The CDF of the retrieval count and retrieval size of posts over an edge is illustrated in Figure 8a and 8b, respectively. For comparison purposes, we also plot the “optimal” case. This assumes that each post is retrieved only once (i.e., from a single relay). We calculate the wasted volume of traffic by subtracting the optimal value from the estimated value. We refer to this as the potentially wasted volume, which is also plotted in the figures.
From Figure 8a and 8b, we observe that the estimated download traffic exchanged is significantly higher than the optimal value (where each post on an edge is retrieved only once). As a result, the estimated amount of traffic wastage is almost identical to the estimated volume of traffic, i.e., nearly all download traffic in Nostr is redundant. On average, the optimal retrieval count for an edge is 1130, while the estimated value is 62839.
To clearly depict this, Figure 8c plots the CDF of the percentage of potentially wasted traffic per follower edge. This graph clearly demonstrates a significant volume of wasted traffic. We observe that 98% of the traffic is wasted for over 50% of the following edges. We find that 98.2% of the download traffic in Nostr is wasted, totalling 144 TiB (Line 12 in Algorithm 1). This is because the majority of clients are connected to multiple relays, thereby redundantly downloading the same objects many times.
Overall, we find that the duplicated retrievals from relays results in substantial traffic waste. This leads to unnecessary costs in operating the relays and potentially has a negative impact on the Nostr ecosystem. It also has a detrimental impact on clients with limited resources. We argue that certain simple innovations could mitigate this issue; thus, we explore possible solutions in §6.3.
@@ 5.4.2. Results
Figure 8. (a) CDF of the retrieval count of posts transmitted over a following edge; (b) CDF of the retrieval size of posts transmitted over a following edge; (c) CDF of the percentage of potentially wasted retrieval count and retrieval size of posts transmitted over a following edge.
@@
inline,color=whiteinline,color=whitetodo: inline,color=whiteTakehomes: (i) The post availability on Nostr significantly surpasses that of Fediverse applications. (ii) However, the overhead of post replications is large. On average, a post is replicated across 34.6 relays, introducing great storage overhead. (iii) Moreover, due to the Nostr protocol and extensive post replications, users retrieve the same post multiple times from different relays, leading to significant wastage of bandwidth for both relays and clients.
6. Design Innovation
In §5, we have demonstrated that the substantial replication of posts in Nostr leads to significant overhead. Consequently, in this section, we investigate ways to enhance it and mitigate the associated overhead. We propose two design improvements: one aimed at decreasing the number of post replications and another focused on minimizing the traffic wasted in post retrieval.
6.1. Improving Storage Redundancy with Strategic Replication Reduction
First, we propose an enhancement for Nostr clients that efficiently constrains the number of post replications. Our goal is to reduce the storage redundancy, while not negatively impacting availability. We refer to our approach as strategic replication reduction. This is motivated by the result in §5.3, which shows that there is significant redundant and potentially unnecessary replication.
In the vanilla Nostr implementation, the client sends the user’s post to each relay recorded in the user’s relay list. This procedure is outlined in pseudo code in Algorithm 2. This, however, results in many redundant replicas. Thus, we propose a simple refinement where, instead of sending to all relays, a function is introduced to designate $N$ relays. Then, the post is exclusively sent to these $N$ relays. This optimized process is outlined in pseudo code in Algorithm 3. Here, we strive to make the relay selection mechanism flexible. The strategy for selecting relays and the number of $N$ should be configurable by the user to suit their specific needs. For the purposes of our evaluation, we next evaluate our proposed strategic replication reduction method using three typical relay selection functions (random, based on AS, and based on user’s followers) with different values of $N$. We will describe their operation in the next section.
6.2. Evaluating Strategic Replication Reduction
We next evaluate our approach. Naturally, a reduction in replication may negatively impact availability. Thus, we aim to (i) evaluate whether, for a reasonable value of $N$ replications, reducing post replications affects post availability; and (ii) provide the client and users a guideline for configuring their own reasonable $N$ value and relay selection functions.
@@ 6.2. Evaluating Strategic Replication Reduction
Evaluation Method.  To assess post availability after limiting replications to a maximum of $N$, we simulate new scenarios with different $N$, derived from our existing dataset. Within the original dataset, each post, $P$, is associated with a map $P\rightarrow S$, where $S$ represents the set of relays where $P$ is replicated. To reflect the scenario where the maximum replications is limited to $N$, this mapping is then modified. For any post $P$ with replications exceeding $N$ (i.e., $|S|>N$), we create a subset $S^{\prime}$ from $S$ such that $|S^{\prime}|=N$, and let $P\rightarrow S^{\prime}$. Subsequently, we employ the same approach as detailed in Section §5.2 to measure the post availability within these new scenarios.
As stated above, we explore three different strategies (relay selection functions) to select the $N$ relays to establish the subset $S^{\prime}$: random, based on ASes, and based on followers. We describe these in detail in the following paragraphs.
Figure 9. Percentage of available posts when top N relays are removed and when top N ASes are removed, if the maximum number of post replications is reduced to N. (a) The reduction is random; (b) The reduction is based on AS; (c) The reduction is based on the followers of the author of the post.
Random Relay Selection.  We begin by measuring the post availability after reducing replications with random relay selection. This naive approach can serve as a baseline. We run the experiments in 4 scenarios with $N=$ 5, 10, 15, and 20, respectively. Additionally, we include the scenario where $N=2$ to illustrate how availability diminishes when the replication number is extremely low. Specifically, for every $P\rightarrow S$ with $|S|>N$, we construct subset $S^{\prime}$ by randomly selecting $N$ relays from $S$, and let $P\rightarrow S^{\prime}$.
The availability results are depicted in Figures 9a. The figure shows that when the maximum post replication is limited to $N\geq 10$, the post availability (percentage of available posts) is actually very close to the original.
To quantify the difference in availability between the new scenario and the original, we begin by defining $A(N,X)$ as the availability when the maximum replication is restricted to $N$, and the top $X$ relays are removed. This corresponds to a specific data point in the figure. Subsequently, we introduce $\Delta^{A}(N,X)=A(N,X)-A(O,X)$, where $O$ denotes the original scenario. This value represents the difference between a given data point and the original. In order to encapsulate the overall difference of a line in the figure and the original, we define $D^{A}_{relay}(N,X)=\Sigma^{X}_{i=1}\Delta^{A}(N,i)/X$. This captures the average difference in availability across all $X$ values.
Using this metric, we calculate $D^{A}_{relay}(20,50)$, $D^{A}_{relay}(15,50)$, and $D^{A}_{relay}(10,50)$ to be -0.05%, -0.18%, and -1.15%, respectively. These values indicate only a marginal decrease in average availability (while the overhead of 380 million to 480 million post replications are removed). This affirmation underscores that the post-availability in these scenarios is comparable to that of the original.
In the case of removing the top $X$ ASes, the figure indicates that the post availability appears satisfactory when the maximum post replications are limited to $N\geq 15$. To quantify this observation, similar to $D^{A}_{relay}$, we define and compute $D^{A}_{AS}(N,X)$ for the removal of $X$ ASes. Upon calculation, we determine that $D^{A}_{AS}(20,30)$ is -2.65%, and $D^{A}_{AS}(15,30)$ is -4.59%. This confirmation underscores that the post availability in these scenarios is comparable to the original.
Overall, the results show that, for reasonably large values of $N$ (such as 15 and 20), reducing the replications with random relay selection does not significantly affect post availability.

### v2

@@ 5. Post Availability & Replication
In §4.3, we find that certain relays (including the larger ones hosting over 1 million posts) experience periods of downtime. However, one of the key design innovations of Nostr is that the posts can be replicated across multiple relays, so that the availability of posts will not be significantly impacted by censorship or failures on a few relays. Indeed, in §4.1 and §4.2, we find many posts are replicated across multiple relays, ASes, and countries. Thus, we next evaluate the effectiveness of this replication design innovation in enhancing post availability and censorship resistance, alongside the overhead associated.
@@ 5.1. Number of Post Replications
We first assess the replication of posts, at a per-post level. Figure 6 presents the CDF of the number of relays and ASes in which a post is replicated. We see remarkable levels of replication: the majority of posts are replicated across multiple relays and ASes, with the average being 34.6 and 12.4, respectively. While the high volume of post replications contributes to increased post availability and censorship-resistance, it also leads to an excess of redundancies. This may incur unnecessary costs, e.g., relay storage space. We delve deeper into this aspect in the subsequent subsections.
5.2. Post Availability under Attack
To evaluate the robustness of posts to censorship attacks, we simulate relay removals using our dataset. Failures could be triggered by censorship attacks from third parties (e.g., DDoS) or simply a shutdown by the administrator. We sequentially remove the most pivotal relays and ASes, and then measure the number of posts that remain available. We employ two ways of ranking the most pivotal relays and ASes to remove: measured by the largest number of (i) posts and (ii) users. We perform an additional ranking for ASes based on the number of relays. Note, we define a post as available if it is replicated on at least one relay.
Figure 7a shows the percentage of available posts after the removal of the top $X$ relays. In stark contrast to the Fediverse (Raman et al., 2019), Nostr’s availability and censorship resistance remains impressively robust, hovering around 90% even after the top 50 relays, ranked by the number of users, are removed. Even when removing the top relays based on the number of posts, the availability remains above 90% only until the top 30th relay is removed (after which it experiences a sharp decline, culminating at 71% at the top 50th). That said, we argue that the availability is still fairly impressive. Figure 7b also displays the proportion of accessible posts following the elimination of the top X ASes. Despite the failure of entire ASes, over 80% of posts remaining available even after the removal of the top 10 ASes. To provide a comparison, the post availability plunges to less than 10% after the removal of the top 10 ASes hosting the instances (Raman et al., 2019). This is largely driven by the more uniform distribution of relays across distinct ASes, making them more resilient to individual AS takedowns. These findings confirm significant censorship resistance within the Nostr ecosystem, a result largely attributable to the extensive replication of posts.
5.3. Post Redundancy
Despite the impressive post availability attained by Nostr, we suspect that this comes at the cost of excessive overhead. This involves an “over-replication” of some posts across numerous (unnecessary) relays. For example, our data shows that, 50% of posts are replicated across over 14 relays, and 25% are replicated across more than 49 relays. Arguably, this level of redundancy is unnecessary, except in the face of catastrophic wide area attacks or failures. Consequently, we next measure the possible optimization space of post replications using our dataset.
The results are depicted in Figure 7c. By restricting the maximum replication to 20, it is possible to eliminate 380 million (61.7%) post replications. Further reducing this limit to 10 leads to the removal of an additional 100 million post replications, totaling 480 million (77.9%). This substantial reduction suggests that, if each user only posts to a maximum of 10-20 relays, over 400 million client-relay communications for posting could be saved, along with the storage space required for the over 400 million replications. We later revisit the implications that this has on resilience to relay failures or censorship takedowns.
@@ 5.4. Duplicated Retrieval From Relays
The prior subsection has highlighted that Nostr’s approach to post replication and substantially improves availability and censorship resistance. However, this comes at the cost of storage overhead. In this subsection, we explore another unexpected overhead created by this approach: A client often must redundantly retrieve the same post from several relays, resulting in unnecessary bandwidth usage. This occurs because in Nostr’s design, a client will maintain connections to all relays the user is using, and retrieve posts from all these relays. For example, if user $A$ follows user $B$, and both are using relays $R_{1}$, $R_{2}$, and $R_{3}$, then $B$ will send a post $P$ to all three relays. Consequently, $A$ will receive the identical post $P$ from $R_{1}$, $R_{2}$, and $R_{3}$ respectively. This is done because a client does not necessarily know which relay will be hosting the next event of interest. However, the ideal scenario would be for $A$ to retrieve $P$ from just one of the three relays. To date, Nostr has no solution to elegantly deal with this duplicated retrieval.
@@
Estimating Traffic Volume.  We begin by estimating the amount of wasted traffic caused by the duplicated retrieval of posts. Here, we assume that a user will retrieve all the posts from the users they are following via all the relays they are using, as defined by the follower relationships. First, we construct a directed graph $G$ of all users to represent following relationships. In this graph, a vertex $V$ signifies a user, and a directed edge $(U,V)$ indicates that user $U$ follows user $V$.
@@
Retrieval Count of a Post over an Edge.  For a post, $P$, that is shared over a following edge $(U,V)$, we calculate the number of times it is retrieved across relays. We refer to this as the retrieval count. First, we get $S_{P}^{R}$, representing the set of relays where the post $P$ is duplicated. Second, we get $S_{U}^{R}$, representing the set of relays utilized by user $U$. Last, we compute $|S_{P}^{R}\cap S_{U}^{R}|$, which corresponds to the retrieval count of $P$.
@@
Traffic over an Edge.  For a following edge $(U,V)$, we calculate the volume of traffic over it. First, we get $S_{V}^{P}$, which represents all posts published by user $V$. Note, these are equal to all posts transmitted over edge $(U,V)$. Second, for every post in $S_{V}^{P}$, we calculate its retrieval count. We then calculate the retrieval size, by multiplying the retrieval count by the size of the post. Last, we sum the retrieval counts and retrieval traffic for all posts within $S_{V}^{P}$ to obtain the total retrieval count and retrieval size over the edge $(U,V)$.
@@
Traffic over the Following Graph.  We also calculate the volume of traffic over the entire follower graph $G$. For every edge in $G$, we compute the traffic over the edge, and sum up the results.
@@
Figure 8. CDF of (a) the retrieval count of posts transmitted over a following edge; (b) the retrieval size of posts transmitted over a following edge; (c) the percentage of wasted retrieval count and retrieval size of posts transmitted over a following edge.
Results.  We use the above method to estimate the quantified amount of traffic. In total, the follower graph $G$ contains 3.5M follower relationships between users. The CDF of the retrieval count and retrieval size of posts over an edge is illustrated in Figure 8a and 8b, respectively. For comparison, we also plot the “optimal” case. This assumes that each post is retrieved only once (i.e., from a single relay). We calculate the wasted volume of traffic by subtracting the optimal value from the estimated value. We refer to this as the wasted volume, which is also plotted in the figures.
From Figure 8a and 8b, we observe that the estimated download traffic exchanged is significantly higher than the optimal value (where each post on an edge is retrieved only once). As a result, the estimated amount of traffic wastage is almost identical to the estimated volume of traffic, i.e., nearly all download traffic in Nostr is redundant. On average, the optimal retrieval count for an edge is 1130 posts, while the actual estimated value is 62,839 posts. This is a significant cost to pay for the additional censorship protections.
To more clearly depict this, Figure 8c plots the CDF of the percentage of wasted traffic per follower edge. This graph shows a significant volume of wasted traffic. We observe that 98% of the traffic is wasted for over 50% of the following edges. We find that 98.2% of the download traffic in Nostr is wasted, totaling 144 TiB . This is because the majority of clients are connected to multiple relays, thereby redundantly downloading the same objects many times.
Overall, we find that the duplicated retrievals from relays results in substantial traffic waste. This leads to unnecessary costs in operating the relays and can have a negative impact on the Nostr ecosystem. It also has a detrimental impact on clients with limited resources. We argue that certain simple innovations could mitigate this issue; thus, we explore possible solutions in §6.2.
@@
inline,color=whiteinline,color=whitetodo: inline,color=whiteTakeaways: (i) Post availability on Nostr significantly surpasses that of Fediverse applications, making censorship difficult without a concerted large-scale attack upon many relays. (ii) However, the overhead of post replications is large. On average, a post is replicated across 34.6 relays, introducing significant storage overhead. (iii) Due to the Nostr protocol and extensive post replications, users retrieve the same post multiple times from different relays, leading to significant wastage of bandwidth for relays and clients.
6. Potential Nostr System Improvements
In §5, we found that the replication of posts in Nostr leads to significant overhead. Thus, we next investigate potential improvement that can reduce the associated overhead. The goal is to provide references and insights for the community to further improve the operation of Nostr, while retaining the same level of censorship resistance.
6.1. Improving Storage Redundancy
§5.3 shows that, by restricting the maximum number of replications, it is possible to eliminate millions of replications. Thus, we propose an enhancement for Nostr clients, where they set a cap of $N$ relays to publish their posts to. This, however, may impact post availability and censorship resistance. We are aware that reducing this redundancy, even slightly, creates a trade-off between theoretical availability and practical efficiency. Our position is that the current level of replication is excessive, imposing a severe burden on relay operators for a level of resilience that may not be necessary in practice. Our goal is to find a more pragmatic balance that ensures the network’s long-term health.
Thus, it is important that a client intelligently selects the most resilient composition of $N$ relays. We strive to make the relay selection mechanism flexible. The strategy for selecting relays and the number of $N$ is therefore configurable by the user to suit their specific needs. Yet, intuitively, we would want a post to be replicated across different ASes to prevent central points of failure. Thus, we propose an AS-based strategy that prioritizes availability, aiming to distribute the post across the maximum possible number of ASes. We evaluate our proposed enhancement by simulating the situation where we limit the maximum number of replicas to $N$ and then compare the post availability to the original using the same method as in §5.2.
@@ 6.1. Improving Storage Redundancy
Evaluation Method.  To assess post availability after limiting replicas to a maximum of $N$, we simulate new scenarios with different $N$ values, derived from our existing dataset. Within the original dataset, each post, $P$, is associated with a map $P\rightarrow S$, where $S$ represents the set of relays where $P$ is replicated. To reflect the scenario where the maximum number of replicas is limited to $N$, this mapping is then modified. For any post $P$ with replicas exceeding $N$ (i.e., $|S|>N$), we create a subset $S^{\prime}$ from $S$ such that $|S^{\prime}|=N$, and let $P\rightarrow S^{\prime}$. To evaluate the trade-off of $N$ and the availability, we employ the same approach as detailed in Section §5.2 (i.e., simulating relay failures) to measure the post availability within these new scenarios. We run the experiments in 4 scenarios with $N=$ 5, 10, 15, and 20, respectively. Additionally, we include the scenario where $N=2$ to illustrate how availability diminishes when the replication number is extremely low.